# PathoDuet: Foundation Models for Pathological Slide Analysis of H&E and IHC Stains

### v1

@@ 4.3 Downstream tasks with IHC images
As for the other part of PathoDuet, to evaluate the IHC model’s competence, we conduct two additional experiments that are closely related to real-world diagnosis, expression level assessment of certain markers, and cancer cell identification.
Assessing the expression level of IHC markers is a primary work for pathologists to assess an IHC slide. We formulate this task as a classification between IHC patches of different expression levels to test the basic capability of models’ tackling IHC images.
Cancer cell identification is also critical in the analysis of IHC images because only the expressed cancer cells are of diagnostic significance. It is yet hard for pathologists to find these cells in IHC slides alone since the structure of single cells is too vague to be recognized. Hence, this task serves as an advanced challenge to models’ understanding of IHC images. Meanwhile, with data from two sites available, this task is intended as a cross-site one to better evaluate models’ generalizing ability.
To note, these tasks are implemented under linear protocol to focus on the inherent capabilities of models. A detailed description is also delivered in the following subsections.
@@ 4.3.1 IHC expression levels assessment
Besides tumor cell identification, assessment of certain marker’s expression levels plays a primary role in IHC diagnosis. We formulate it as a patch-level multi-class classification task, because trivial regression may lose focus on certain scores that have diagnostic meaning. We manually selected several thresholds closely related to pathologists’ examination. A linear probing method is applied with metrics including Acc, balanced accuracy (bAcc), and weighted F1 (wF1) score because of class imbalance.
@@
Results. In Table 4, the performance of different models is reported in different amounts of training data. Reviewing the overall results, we can suppose that pathological foundation models of H&E images can have a certain insight into IHC images because we can see superior performance over ImageNet models. When focusing on individual models, our IHC model demonstrates attracting performance on most metrics, especially in the limited training data case. Notably, CTransPath also presents excellent performance, especially when training data is sufficient. This can be credited to the superiority of fusing CNN and Swin Transformer, as well as the usage of PAIP dataset  [^2] http://wisepaip.org/paip, which contains some IHC slides. When the difficulty of the task increases, the advantage of explicit transferring to the IHC model is more obvious. The phenomenon states that although H&E models can provide satisfactory results with adequate training data, explicit transferring is necessary for limited annotated data.
4.3.2 Cross-site tumor identification
Datasets. The dataset is also private, consisting of IHC images from two medical sites. The annotation is simply positive (images containing tumor cells) and negative (images without any tumor cells). The first site contains two sets of data stained with ER, KI67, and PR acquired at different times, resulting in slightly different appearances. The first (Site 1) has 1,365 patches (700 positive and 665 negative), and the other has 126 patches (64 positive and 62 negative), denoted as Set L and Set S respectively. In the second site (Site 2), there are 3,688 patches with much more domain shift, including 2,684 positive and 1,004 negative stained with ER, KI67, and EGFR.
To better evaluate the models’ performance, we design four scenarios. Set L normal-$\rightarrow$ Set S and Set S normal-$\rightarrow$ Set L are viewed as in-site settings. The difference is we use a larger dataset (Set L) as the training set in the former one and a smaller one in the latter one. Data in Site 2 can be viewed as being sampled from a new domain other than Site 1. To this, it can be used for out-of-distribution (OOD) evaluation. We also design two OOD settings. Site 2 Seen keeps 20% of data in Site 2, and uses it for model selection. The model is trained with Set L in Site 1, validated with 20% data in Site 2, and tested with the rest 80% data. Site 2 unseen is a stricter setting, which is also common in real-world scenarios. The model is trained with Set L in site 1, validated with Set S, and tested with data from absolutely unseen Site 2.
@@ 4.3.2 Cross-site tumor identification
Results. In Table 5, the performance of different models is reported under four evaluation scenarios. When applying in-site settings, ImageNet supervised ViT, CTransPath, and our IHC model all present remarkable performance. When tested on another site, ImageNet models yet display a weakening performance. Meanwhile, our model offers the best performance in a setting of partially available data from another domain and maintains the performance in a pure OOD setting. Other pathological models perform well with some a priori information from another domain, but may not keep it when the test set is totally unseen. Reviewing the overall results, we can derive similar conclusions. Foundation models of H&E images can be a good choice to understand IHC images, compared with models trained with natural images. An IHC-specific model, however, can provide more insights especially when the difficulty of the task is relatively high, i.e., large domain shift and limited annotated data.
@@ 5 Discussion
In this section, we first conduct several ablation experiments to validate the efficacy of two proposed pretext tasks. Then, we study the data requirement of applying foundation models to downstream tasks, to demonstrate the necessity of developing pathological foundation models. Finally, we compare our H&E model to some giant model pretrained with ultra-large scale pathological datasets and observe that both increasing the amount of data and designing a tailored framework help to develop powerful pathological models.
5.1 Efficacy of the proposed methods
In Section 3.2, we claim that the essence of the cross-scale positioning task is to bridge representations in a local view and those in a global view. To validate whether cross-scale positioning enhances models’ global understanding, we choose the WSI classification task since it is more related to a global understanding of a slide. We compare the performance of a purely MoCo v3 pretrained model using our dataset and the model using cross-scale positioning, named “MoCo v3” and “+XSP”, respectively. From the results in Table 6, we can see that the later model outperforms the purely MoCo v3 one in most metrics. The results can prove that cross-scale positioning boosts models’ understanding of an image from a broader view. Besides, cross-scale positioning also improves overall performance, as shown in Table 7.
@@ 5.2 Data requirements of downstream tasks
Figure 3: Data requirement study on NCT-CRC-HE dataset. PathoDuet is compared with ImageSup and the performance with the full dataset as an upper bound is represented by the dotted line with the same color.
In Fig. 3, we compare our H&E model with ImageNet-supervised ViT using the NCT-CRC-HE dataset. We can observe that although these two models can present close performance with the full dataset ($10^{6}$ images) when the amount is limited to only $10^{2}$, the performance gap is large. This can be used to validate the reduced data requirement when using pathological foundation models.
Figure 4: Data requirement study on different datasets using PathoDuet.
@@ 5.3 Comparison with giant pathological models
Recently, there has been a trend in pretraining pathological models with astronomical numbers of diagnostic slides. UNI, proposed by Chen et al. (2023), utilizes 100,000 slides to train a ViT-Large with DINO v2 framework. Virchow (Vorontsov et al., 2023) further extends this number to 1.5 million and employs a ViT-Huge model with DINO v2 as well. In this section, we compare our model to those giants and CTransPath as a baseline in the CRC patch classification task. Notably, except our models, all the figures are copied from Virchow’s experiment, so the results of CTransPath are different from that in Section 4.2.1, which is as small as 0.002 in accuracy, and 0.001 in balanced accuracy and weight F1 score, and can be owed to different training strategies and randomness. Here, we provide two versions of our model, a normal ViT used throughout the aforementioned experiments, and a ViT with our positioner network to aggregate the features instead of using the vanilla average. The latter one is marked with a *.
@@ 6 Conclusion
We introduce PathoDuet, a series of foundation models on computational pathology, covering both H&E and IHC images, and propose a new self-supervised learning framework with two pretext tasks in pathology. The key to this framework is the introduction of a pretext token and following task raisers. It consists of both a model pretraining task, cross-scale positioning, and a model adaptation task, cross-stain transferring. In cross-scale positioning, we bridge the local and global representations of H&E patches to enhance pathological image understanding in various magnifications. In cross-stain transferring, we utilize adaptive instance normalized H&E features to provide pseudo-IHC features injected with structural information. The original H&E model is therefore transferred to an interpreter of IHC images. We evaluate the performance of our models over a wide variety of downstream tasks, and the experimental results show the efficacy of our models on most tasks. Besides, we also investigate the downstream data requirements and comparison with giant pathological models, to discover the power of data and delicately designed SSL methods tailored to pathological images.
Supplementary Material
Detailed Experimental Settings. For linear evaluation on the NCT-CRC-HE dataset, following CTransPath, we use Adam as the optimizer with a batch size of $96$. The initial learning rate is set to $0.0003$. Data augmentations of random horizontal, vertical, and 90-degree flips and random cropping are used. For the full fine-tuning, the learning rate for the linear classifier is set to $0.0003$, while the learning rate for the rest part of the pretrained model is set to $0.00003$. The maximum number of epochs is set to $50$ for all models to converge.
For the weakly-supervised WSI classification task, following CLAM, we freeze our pretrained model and use the Adam as the optimizer with a batch size of 1 (WSI/bag) and a weight decay of $0.00001$. The learning rate is set to $0.0002$, and the epochs as $50$. For the CAMELYON16 dataset, we adopt the official data split in the CAMELYON16 challenge. For TCGA-NSCLC and TCGA-RCC datasets, we use 5-fold Monte Carlo cross-validation to obtain a more stable result. Each WSI is cropped into non-overlapping small patches after removing the background area with a filter of saturation less than 15 in the CAMELYON16 dataset, and 8 in the TCGA dataset.
For IHC expression levels assessment, we also keep the linear classification setting. The batch size is $128$ for full training data and $64$ for 5% training data. The learning rate is $0.02$. For IHC tumor identification, a linear probing method is implemented as well. In Set L normal-$\to$ Set S, Site 2 Seen and Site 2 Unseen, the batch size is $128$ and the learning rate is $0.005$. In Set S normal-$\to$ Set L, the batch size is reduced to $64$. The maximum number of epochs is also set to $50$ for convergence of all models.

### v2

@@ 4.3 Downstream tasks with IHC images
As for the other part of PathoDuet, to evaluate the IHC model’s competence, we conduct three additional experiments that are closely related to real-world diagnosis, expression level assessment of certain markers, cancer cell identification and slide-level IHC qualitative analysis.
Assessing the expression level of IHC markers is a primary work for pathologists to assess an IHC slide. We formulate this task as a classification between IHC patches of different expression levels to test models’ basic capability of tackling IHC images. Cancer cell identification is also critical in the analysis of IHC images because only the expressed cancer cells are of diagnostic significance. It is yet hard for pathologists to find these cells in IHC slides alone since the structure of single cells is too vague to be recognized. Hence, this task serves as an advanced challenge to models’ understanding of IHC images. Meanwhile, with data from two sites available, this task is intended as a cross-site one to better evaluate models’ generalizing ability. To note, these tasks are implemented under linear protocol to focus on the inherent capabilities of models. Slide-level IHC qualitative analysis tests models’ performance on WSI diagnosis. We collect three different markers, and exploit multiple instance learning method to ask models to give a positive or negative prediction of certain marker. A detailed description is also delivered in the following subsections.
@@ 4.3.1 IHC expression levels assessment
Besides tumor cell identification, assessment of certain marker’s expression levels plays a primary role in IHC diagnosis. We formulate it as a patch-level multi-class classification task, because trivial regression may lose focus on certain scores that have diagnostic meaning. We manually selected several thresholds closely related to pathologists’ examination. A linear probing method is applied with metrics including accuracy (Acc), balanced accuracy (bAcc), and weighted F1 (wF1) score because of class imbalance.
@@
Settings. For IHC expression levels assessment, we also keep the linear classification setting in Section 4.2.1. The batch size is $128$ for full training data and $64$ for 5% training data. The learning rate is $0.02$.
Results. In Table 4, the performance of different models is reported in different amounts of training data. Reviewing the overall results, we can suppose that pathological foundation models of H&E images can have a certain insight into IHC images because we can see superior performance over ImageNet models. When focusing on individual models, our IHC model demonstrates attracting performance on most metrics, especially in the limited training data case. Notably, CTransPath also presents excellent performance, especially when training data is sufficient, and SimCLR-ciga provides second best performance when training data is limited. This might be credited to their use of some IHC images as pretraining data. Besides, although UNI does not see IHC images during pretraining, it still presents surprisingly powerful generalizing ability to IHC images. When the difficulty of the task increases, the advantage of explicit transferring to the IHC model is more obvious. The phenomenon states that although H&E models can provide satisfactory results with adequate training data, explicit transferring is necessary for limited annotated data.
@@
Table 6: Results of slide-level prediction of CD5. The best performance in each column is bold, and the second best is underlined.
Methods AUC Acc F1 Recall Precision Specificity
ImageSSL 0.931 0.888 0.908 0.909 0.908 0.860
ImageSup 0.924 0.873 0.895 0.895 0.899 0.844
SimCLR-ciga 0.862 0.798 0.837 0.850 0.826 0.722
RetCCL 0.881 0.825 0.852 0.840 0.866 0.803
CTransPath 0.868 0.814 0.844 0.834 0.856 0.784
UNI 0.907 0.863 0.887 0.888 0.887 0.828
Ours (IHC) 0.924 0.840 0.862 0.835 0.895 0.850
4.3.2 Cross-site tumor identification
Datasets. The dataset is also private, consisting of IHC images from two medical sites. The annotation is simply positive (images containing tumor cells) and negative (images without any tumor cells). The first site (Site 1) contains two sets of data stained with ER, KI67, and PR acquired at different times, resulting in slightly different appearances. The first set has 1,365 patches (700 positive and 665 negative), and the other has 126 patches (64 positive and 62 negative), denoted as Set L and Set S respectively. In the second site (Site 2), there are 3,688 patches with much more domain shift, including 2,684 positive and 1,004 negative stained with ER, KI67, and EGFR.
To better evaluate the models’ performance, we design four scenarios. Set L $\rightarrow$ Set S and Set S $\rightarrow$ Set L are viewed as in-site settings. The difference is we use a larger dataset (Set L) as the training set in the former one and a smaller dataset (Set S) in the latter one. Data in Site 2 can be viewed as being sampled from a new domain other than Site 1. To this, it can be used for out-of-distribution (OOD) evaluation. We also design two OOD settings. Site 2 Seen keeps 20% of data in Site 2, and uses it for model selection. The model is trained with Set L in Site 1, validated with 20% data in Site 2, and tested with the rest 80% data. Site 2 unseen is a stricter setting, which is also common in real-world scenarios. The model is trained with Set L in site 1, validated with Set S, and tested with data from absolutely unseen Site 2.
Settings. For IHC tumor identification, a linear probing method is implemented as well. In Set L $\to$ Set S, Site 2 Seen and Site 2 Unseen, the batch size is $128$ and the learning rate is $0.005$. In Set S $\to$ Set L, the batch size is reduced to $64$. The maximum number of epochs is also set to $50$ for convergence of all models.
@@ 4.3.2 Cross-site tumor identification
Results. In Table 5, the performance of different models is reported under four evaluation scenarios. When applying in-site settings, ImageNet supervised ViT, CTransPath, UNI, and our IHC model all present remarkable performance. When tested on another site, ImageNet models yet display a weakening performance. Meanwhile, our model offers the best performance in a setting of partially available data from a new domain and maintains the performance in a pure OOD setting. Other pathological models perform well with some a priori information from another domain, but may not keep it when the test set is totally unseen.
Table 7: Results of slide-level prediction of CD10. The best performance in each column is bold, and the second best is underlined.
Methods AUC Acc F1 Recall Precision Specificity
ImageSSL 0.898 0.842 0.813 0.802 0.827 0.873
ImageSup 0.870 0.800 0.758 0.737 0.784 0.848
SimCLR-ciga 0.902 0.842 0.810 0.814 0.824 0.874
RetCCL 0.855 0.782 0.742 0.749 0.752 0.819
CTransPath 0.849 0.773 0.729 0.720 0.754 0.825
UNI 0.912 0.828 0.786 0.757 0.820 0.880
Ours (IHC) 0.909 0.835 0.795 0.766 0.838 0.894
Table 8: Results of slide-level prediction of CD21. The best performance in each column is bold, and the second best is underlined.
Methods AUC Acc F1 Recall Precision Specificity
ImageSSL 0.677 0.614 0.667 0.690 0.662 0.535
ImageSup 0.638 0.571 0.614 0.628 0.631 0.520
SimCLR-ciga 0.659 0.582 0.638 0.660 0.627 0.492
RetCCL 0.657 0.604 0.659 0.688 0.644 0.507
CTransPath 0.701 0.637 0.667 0.651 0.703 0.631
UNI 0.635 0.607 0.644 0.641 0.658 0.570
Ours (IHC) 0.740 0.626 0.669 0.695 0.667 0.548
4.3.3 Qualitative analysis of IHC slides
Besides patch-level IHC tasks, diagnosis directly from IHC slides is also of great importance. Therefore, we collect some IHC slides of different markers, and invite some experts to give a positive or negative label to each slide. This task further examines models’ capability of assessing IHC slides given certain marker. We still apply CLAM as the training method. The evaluating metrics include AUC, Acc, F1 score, Recall, Precision and Specificity.
Datasets. The dataset includes 3 IHC markers, CD5, CD10 and CD21. For each marker, we collect over 250 slides and annotate each slide with a positive or negative label. To be detailed, we collect 124/189 positive/negative CD5 slides, 139/111 CD10 slides, and 115/150 CD21 slides, respectively. We use each marker solely and implement 5-fold cross validation.
Settings. For slide-level prediction of IHC markers, we follow the settings used in TCGA experiments as introduced in Section 4.2.2.
Results. In Table 6, 7 and 8, we present the results of slide-level prediction of CD5, CD10 and CD21. If we take three tables as a whole, we can see that ImageSSL, SimCLR-ciga, CTransPath, UNI and our models provide relatively good and stable performance. For ImageSSL, an interesting fact is that it surpasses ImageSup in all statistics. A reasonable guess is self-supervised learning helps model generalizing to other domains, especially those rarely-seen domains. For SimCLR-ciga and CTransPath, as aforementioned, they may benefit from some IHC data used for pretraining. For UNI, it still presents great performance with its surprising generalizing ability. For our model, it provides outstanding performance, achieving at least second-best AUC across all three datasets, while no other models can achieve. This should be credited to cross-stain transferring, which exploits existing H&E model and a little amount of paired IHC and H&E data. Reviewing the overall results, we can derive similar conclusions. Foundation models of H&E images can be a good choice to understand IHC images, compared with models trained with natural images. An IHC-specific model, however, can provide more insights especially when the difficulty of the task is relatively high, i.e., large domain shift and limited annotated data.
@@ 5 Discussion
In this section, we first conduct several ablation experiments to validate the efficacy of two proposed pretext tasks. Then, we study the data requirement of applying foundation models to downstream tasks, to demonstrate the necessity of developing pathological foundation models. Finally, we compare our H&E model to some giant models pretrained with ultra-large scale pathological datasets and observe that both increasing the amount of data and designing a tailored framework help to develop powerful pathological models.
@@
5.1 Efficacy of the proposed methods
In Section 3.2, we claim that the essence of the cross-scale positioning task is to bridge representations in a local view and those in a global view. To validate whether cross-scale positioning enhances models’ global understanding, we choose the WSI classification task since it is more related to a global understanding of a slide. We compare the performance of a purely MoCo v3 pretrained model using our dataset and the model using cross-scale positioning, named MoCo v3 and +XSP, respectively. From the results in Table 9, we can see that the latter model outperforms the purely MoCo v3 one in most metrics. The results can prove that cross-scale positioning boosts models’ understanding of an image from a broader view. Besides, we also conduct ablating experiments on NCT-CRC-HE dataset, and observe a little performance gain after applying cross-scale positioning task as shown in Table 10. The margin in full finetuning setting barely exists, which is reasonable since the two models share the same model architecture. If we take the results in Table 9 and Table 10 as a whole, we can derive the conclusion that cross-scale positioning helps model understand H&E images better.
@@ 5.2 Data requirements of downstream tasks
Fig. 4: Data requirement study on NCT-CRC-HE dataset. PathoDuet is compared with ImageSup and the performance with the full dataset as an upper bound is represented by the dotted line with the same color.
In Fig. 4, we compare our H&E model with ImageNet-supervised ViT using the NCT-CRC-HE dataset. We can observe that although these two models can present close performance with the full dataset ($10^{6}$ images), when the amount is limited to only $10^{2}$, the performance gap is large. This can be used to validate that using pathological foundation models can reduce data requirement of downstream tasks in pathology.
Fig. 5: Data requirement study on different datasets using PathoDuet.
@@ 5.3 Comparison with giant pathological models
Recently, there has been a trend in pretraining pathological models with astronomical numbers of diagnostic slides. UNI, proposed by Chen et al. [6], utilizes 100,000 slides to train a ViT-Large with DINO v2 framework. Virchow [49] further extends this number to 1.5 million and employs a ViT-Huge model with DINO v2 as well. In this section, we compare our model to those giants and CTransPath as a baseline in the CRC patch classification task. Notably, except our models, all the figures are copied from Virchow’s experiment, so the results of CTransPath are different from that in Section 4.2.1, which is as small as 0.002 in accuracy, and 0.001 in balanced accuracy and weight F1 score, and can be owed to different training strategies and randomness. Here, we provide two versions of our model, a normal ViT used throughout the aforementioned experiments, and a ViT with our positioner network to aggregate the features instead of using the vanilla average. The latter one is marked with a *.
@@ 6 Conclusion
We introduce PathoDuet, a series of foundation models on computational pathology, covering both H&E and IHC images, and propose a new self-supervised learning framework with two pretext tasks in pathology. The key to this framework is the introduction of a pretext token and following task raisers. It consists of both a model pretraining task, cross-scale positioning, and a model adaptation task, cross-stain transferring. In cross-scale positioning, we bridge the local and global representations of H&E patches to enhance pathological image understanding in various magnifications. In cross-stain transferring, we utilize adaptive instance normalized H&E features to provide pseudo-IHC features injected with structural information. The original H&E model is therefore transferred to an interpreter of IHC images. We evaluate the performance of our models over a wide variety of downstream tasks, and the experimental results show the efficacy of our models on most tasks. Besides, we also investigate the downstream data requirements and comparison with giant pathological models, to discover the power of data and delicately designed SSL methods tailored to pathological images. PathoDuet highlights the importance of training strategy, while the giants, UNI and Virchow, point out the advantage of preparing sufficient training data. Hence, we will take all efforts to collect more data to iterate and upgrade our models in the future.
Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62301311).